### v1

# Transaction Profiling and Address Role Inference in Tokenized U.S. Treasuries

Abstract
Tokenized U.S. Treasuries have emerged as a prominent subclass of real-world assets (RWAs), offering cryptographically enforced, yield-bearing instruments collateralized by sovereign debt and deployed across multiple blockchain networks. While the market has expanded rapidly, empirical analyses of transaction-level behaviour remain limited. This paper conducts a quantitative, function-level dissection of U.S. Treasury-backed RWA tokens including BUIDL, BENJI, and USDY, across multi-chain: mostly Ethereum and Layer-2s. We analyze decoded contract calls to isolate core functional primitives such as issuance, redemption, transfer, and bridge activity, revealing segmentation in behaviour between institutional actors and retail users. To model address-level economic roles, we introduce a curvature-aware representation learning framework using Poincaré embeddings and liquidity-based graph features. Our method outperforms baseline models on our RWA Treasury dataset in role inference and generalizes to downstream tasks such as anomaly detection and wallet classification in broader blockchain transaction networks. These findings provide a structured understanding of functional heterogeneity and participant roles in tokenized Treasury in a transaction-level perspective, contributing new empirical evidence to the study of on-chain financialization.
keywords:
tokenized U.S. Treasuries, address role inference, transaction function analysis
category:
[…]
1 Introduction
Real-world assets (RWAs) refer to off-chain, yield-bearing financial instruments such as sovereign debt, treasury, credit products, and real estate that are instantiated on-chain as cryptographically enforceable claims via tokenized representations [12], typically underpinned by compliance-bound issuance frameworks and asset custody infrastructures [21]. Early concept and implementation of RWA tokenization in decentralized finance (DeFi) pioneered in 2019 with the launch of Centrifuge’s Tinlake platform [22], which enabled the issuance of ERC-20 tokens [5] backed by tokenized real-world assets such as invoices and trade receivables. Subsequently, RWA tokens evolved to encompass a broader class of fixed-income and yield-generating instruments, notably short-duration U.S. Treasuries [18], tokenized real estate [24], private credit loans [3], each implemented through on-chain implementation of financial contracts governed by legal wrappers, custodial attestations, and programmable functions embedded in ERC-20 [5] or ERC-1400 [23] token standards.
Tokenized representations of short-duration U.S. Treasury bills have emerged as a dominant subclass within the RWA market, constituting over \$6.1 billion in on-chain assets as of Q2 2025 [20]. Leading instruments include BUIDL (BlackRock), OUSG (Ondo Finance), STBT (Matrixdock), and BENJI (Franklin Templeton) [15]. Economically, these tokens function as digital cash equivalents collateralized by Treasury securities, with daily yield accruals implemented via rebasing mechanics (e.g., rOUSG) [14] or dividend distribution (e.g., BUIDL) [25]. Smart contracts encode fund accounting logic, including NAV maintenance, dividend accrual, and investor disbursements, with compliance modules (e.g., Securitize) enforcing identity gating and transfer restrictions. The composability of these instruments within permissionless DeFi remains structurally limited due to non-transferability constraints and KYC-enforced whitelists. Despite this growing RWA U.S. treasury market, the on-chain behavioural patterns of these instruments across issuance, transfer, yield distribution, and institutional wallet interactions remain largely unexamined from a data-driven, motivating transaction-level analysis and network modeling of the emerging Treasury token market.
Prior empirical and smart contract-oriented research on RWA tokenization has centred on structural features of asset issuance and transfer mechanisms [4, 7], highlighting institutional governance trade-offs [28], composability limitations [7], and participation frictions [27, 11], while offering limited analysis of transactional behaviour from a data-driven perspective and cross-chain functional heterogeneity at scale. While those previous efforts were undertaken on RWA tokenization across various asset classes, focused empirical analysis of U.S. Treasury-backed tokens remains limited, even as these instruments have grown significantly since recent years, with increasing contract deployments and transaction data generated flows spanning various financial functionalities.
To address persistent analytic gaps in the study of tokenized real-world assets, this work conducts a function-decoded, cross-chain dissection of tokenized U.S. Treasury transaction networks, isolating transaction-level functional primitives: issuance, redemption, transfer, and bridging, to delineate institutional versus retail financial behaviours. This multi-chain functional analysis enables the differentiation of institutional and retail usage patterns by comparing transaction frequency, notional value, and function-specific activity both within individual chains and in aggregate across the RWA token protocols. The analysis is grounded in transaction-level data collected across multiple chains for three representative U.S. Treasury-backed RWA tokens: BUIDL (BlackRock USD Institutional Digital Liquidity Fund), BENJI (Franklin OnChain U.S. Government Money Fund), and USDY (Ondo U.S. Dollar Yield). We further propose a role classification model for addresses by integrating hyperbolic representation learning, liquidity-based behavioural features, and hierarchical graph descriptors to infer latent economic functions such as treasuries bots, and retail traders.
1.1 Overview
To the best of our knowledge, this is the first study to systematically collect, decode, and structure transaction-level data for tokenized U.S. Treasuries RWA tokens, supporting a quantitative, function-level analysis of these assets from a data-driven perspective. Our contributions are as follows:
  • Cross-chain functional decomposition We present a contract-decoded, function-level analysis of tokenized U.S. Treasuries across multiple chains, revealing how issuance, redemption, transfer, and DeFi-related operations vary by chain and token design.
  • Investor type via transactional profiling: We perform statistical segmentation of wallet behaviour using transaction frequency and value distributions, identifying differences between institutional-scale actors and retail participants. Our results show that institutional wallets dominate issuance and redemption flows, while retail users engage primarily in mid-value transfer activity concentrated on Layer-2 networks.
  • Address role inference via curvature-aware representation learning: We propose a predictive framework based on Poincaré embeddings to capture latent transaction geometry for address role inference (e.g., treasury, bots, traders). Our approach outperforms established baselines on the RWA Treasury dataset and exhibits competitive generalization to downstream tasks such as anomaly detection and entity classification on diverse public blockchain transaction network datasets.
2 Related Work
Real-world asset (RWA) tokenization has increasingly focused on fixed-income instruments, with recent studies examining blockchain-based representations of sovereign debt, particularly tokenized U.S. Treasuries, as a representative use case within on-chain capital markets [4]. A recent multi-sector case study similarly reports that tokenization can improve transaction efficiency and create new value, yet it also introduces governance complexities and shifts risk distribution in on-chain markets [28]. In addition, empirical surveys of thirty-nine RWA projects have been studied to reveal common on-chain vulnerabilities, such as heavy reliance on stablecoins for settlement and a limited base of active on-chain investors [7].
From a behavioural perspective, research efforts have been undertaken examine address-level data to characterize RWA usage patterns. Swinkels studied Ethereum-based real estate tokens (fractional shares of 58 U.S. rental properties) and found highly fragmented ownership, with around 254 unique holders per property on average [27]. The author observed that larger token holders tended to spread investments across multiple properties, while overall liquidity was low: each property token changed ownership only about once per year (slightly more often if tradable on decentralized exchanges). Kreppmeier et al. tracked 173 U.S. real estate security tokens (over 238,000 on-chain transactions) and similarly found broad participation by small investors, though individual wallets were not well-diversified across different tokens [11]. In addition to property fundamentals, the study showed that crypto-market factors such as transaction costs and investor sentiment significantly influence both initial token offering success and subsequent trading flows. These studies demonstrate the value of categorizing on-chain addresses by their roles and behaviours, for instance, distinguishing issuers, custodians, and various types of investors to better interpret RWA economical patterns. The emerging of U.S. Treasury RWA tokens, i.e., on-chain shares in Treasury-backed funds, has drawn interest for bringing safe assets on-chain [18], although academic analysis of their transaction-level semantics such as participant composition remains scant.
3 Data Collection
We collect raw on-chain transactions for tokenized real-world asset products representing U.S. Treasuries, namely, BUIDL (BlackRock USD Institutional Digital Liquidity Fund), BENJI (Franklin OnChain U.S. Government Money Fund), and USDY (Ondo U.S. Dollar Yield) across multiple blockchains using public blockchain explorer APIs. Our collection pipeline queries EVM-compatible chains (e.g., Ethereum, Arbitrum, Mantle), as well as non-EVM chains (e.g., Aptos, Stellar), and fetches transfers for the three most prominent RWA tokens by market capitalization: BUIDL, BENJI, and USDY, based on rankings from market cap  [^1] https://app.rwa.xyz/treasuries.
Table 1: Cross-chain summary statistics for tokenized U.S. Treasury instruments. Columns report the number of transactions, unique participating addresses, and the range of observed activity.
Token Chain Transaction Address Start End
USDY Arbitrum 13,296 1,281 2024-08-12 2025-04-28
Aptos 10,000 2,259 2024-08-31 2025-04-22
Ethereum 2,921 729 2023-09-18 2025-04-21
Mantle 274,657 7,221 2023-12-23 2025-04-22
BENJI Polygon 996 5 2023-10-03 2025-04-23
Arbitrum 1,112 7 2023-11-13 2025-04-25
Base 289 4 2024-11-20 2025-04-23
Avalanche 356 5 2024-10-11 2025-04-23
Aptos 142 3 2024-10-01 2025-04-23
Ethereum 212 3 2024-11-20 2025-04-23
Stellar 2,546,750 220,784 2024-02-22 2024-03-06
BUIDL Ethereum 4,639 65 2024-03-04 2025-04-21
Polygon 204 7 2024-11-04 2025-04-21
Arbitrum 135 6 2024-11-04 2025-04-21
Optimism 136 4 2024-11-04 2025-04-21
Avalanche 25 6 2024-11-04 2024-12-06
BUIDL shows the most activity on Ethereum, aligning with its institutional adoption, though it is collected across multiple chains. The BENJI token exhibits exceptionally high activity on Stellar, with over 2.5 million transactions involving more than 220,000 addresses in just two weeks, indicating high-frequency issuance or custodial activity. USDY shows dense activities on Mantle, accounting for more than 274,000 transactions. Overall, this collected multi-chain dataset enables us to quantify RWA adoption patterns and behavioural clusters in later sections. To decode smart contract method calls, we utilize Tenderly’s decoding engine  [^2] https://tenderly.co/ and the open-source 4byte signature database  [^3] https://www.4byte.directory/ to decode the input field of each transaction, enabling function-level analysis. The result is a cross-chain, function decoded dataset of RWA token transfers suitable for downstream quantitative analysis.
4 Quantitative Analysis of RWA Tokens
To quantitatively assess whether U.S. Treasury token transactions align more closely with institutional investors or retail users, we analyze the frequency of transactions versus their size. An institutional usage pattern is characterized by relatively few transactions, each of very large notional value, while a retail-driven pattern involves a high frequency of smaller transactions, typically ranging from a few dollars to several thousand. We further stratify this analysis by both chains and functions, to assess how usage patterns vary across blockchain environments (e.g., Ethereum vs. Arbitrum) and operational roles (e.g., issuance, redemption, transfer). This enables us to disentangle how token delivers differently in function-level interaction and chain-specific deployment.
Figure 1: Log-scaled scatter plot of BUIDL token interactions by function and chain. Each point represents a distinct (function, chain) pair, where the x-axis is the total BUIDL value transacted and the y-axis is the transaction count. Marker shape denotes functional class (e.g., transfer, issuetokens, redeem), and colour indicates blockchain deployment (e.g., Ethereum, Arbitrum). Institutional behaviour is typically clustered at the top-right (few, large transactions), while retail-aligned activity appears bottom-left (many, small transactions).
4.1 BUIDL
Table 2: Functional buckets for decoded contract functions. Substrings are matched against decoded function names. Function reflect high-level operational roles in RWA token flows.
Function buckets Name Match Description
issuetokens issue, mint, bulkissuance Initial minting and dividend payments to investors
redeem redeem Investor withdrawals
burn burn Permanent destruction of tokens
transfer transfer, bridgedstokens, multisend Routine transfers or bridging between chains
deposit deposit Vault top-ups or liquidity provisioning
deliver deliver Fee or metadata delivery helpers
We categorize smart contract decoded functions into operational buckets based on string-matching heuristics against lowercase function names to interpret the functional semantics. Table 2 summarizes these mappings, which allow for cross-chain, token-agnostic aggregation of function activity into economically meaningful categories. For instance, the issuetokens bucket includes all minting operations ranging from initial token creation to dividend payment (e.g., via batched bulkIssuance calls in BUIDL). Likewise, redeem captures capital outflows by investors, while burn reflects permanent supply contraction. The transfer bucket encompasses routine value transfers, including bridged hops and multi-send operations.
Figure 1 visualizes the distribution of BUIDL token interactions by function and chain, where each data point corresponds to a unique $(f,c)$ pair: a specific function (bucket) $f$ executed on chain $c$. The x-axis (log scale) captures the total value transferred through the function, while the y-axis captures the number of transactions. Marker shapes denote function classes (e.g., transfer, issuetokens), and colours distinguish the underlying chains (e.g., Ethereum, Polygon, Arbitrum). Points located in the upper-right quadrant represent functions with both high frequency and high cumulative value. In contrast, points near the upper-left indicate high-frequency, low-value interactions consistent with retail-like activity, and points in the lower-right denote low-frequency but high-value functions, as expected for capital-intensive institutional transfers.
The observed distribution of BUIDL’s on-chain activity is heavily skewed toward the lower-right, especially on Ethereum, with large-value, low-frequency function calls such as redeem, burn, and issuetokens dominating the volume. This pattern is consistent with usage by accredited or institutional actors, rather than mass-market retail adoption. The dominance of Ethereum reinforces its role as the canonical execution environment for regulated, institutional-grade RWA issuance and settlement. However, the presence of BUIDL transactions on alternative chains such as Avalanche, Polygon, and Optimism indicates emerging interoperability requirements. These deployments may support secondary custody flows, bridging infrastructure, or protocol-level integrations where gas efficiency or modular composability is prioritized. Such cross-chain activity for BUIDL remains limited as of April 2025, based on our collected data, with the vast majority of transaction volume and functional engagement concentrated on Ethereum. Notably, BUIDL transactions are almost exclusively initiation-side operations (e.g., primary issuance and redemption), with minimal secondary trading activity, suggesting that most interacting addresses represent institutional counterparties rather than retail participants.
4.2 BENJI
Franklin Templeton’s BENJI token representing shares of its OnChain U.S. Government Money Fund, was initially deployed on the Stellar blockchain in 2021 and has since used Stellar as the primary ledger for recording share ownership and transactions [26]. In 2023–2024, BENJI was expanded beyond Stellar onto several EVM-compatible chains: Polygon, Arbitrum, Avalanche, and Ethereum Mainnet. Each blockchain in the BENJI token’s plays a specialized role, forming an integrated cross-chain architecture rather than independent silos. Stellar still functions as the primary issuance and settlement network, where the majority of BENJI tokens are minted and held (most of the token supply was on Stellar as of April 2024 [9]).
Table 3: Transaction summary for BENJI across supported chains and decoded functions, based on data collected through April 2025. Stellar shows exceptionally high activity due to its role as the fund’s primary settlement ledger, while EVM chains primarily use the signedDataExecution function for controlled, authorized operations.
Chain Function Transanction Total Value First Seen Last Seen
Arbitrum signeddataexecution 622 192,720,281 2023-11-13 2025-04-25
Avalanche signeddataexecution 353 34,645,861 2024-10-11 2025-04-23
Base signeddataexecution 185 31,521,747 2024-11-20 2025-04-23
Ethereum signeddataexecution 185 30,459,341 2024-11-20 2025-04-23
Polygon signeddataexecution 938 30,640,118 2023-10-03 2025-04-23
Stellar - 2,546,750 - 2024-02-22 2024-03-06
The BENJI token’s on-chain activity demonstrates a stark bifurcation between Stellar and EVM-compatible chains. On Stellar, over 2.5 million transactions were recorded within 14 days, primarily driven by operational activity such as dividend reinvestments, account initializations, and share registry updates. These actions occur at high frequency and low value per transaction, and although their financial values are not directly embedded in the transaction logs, they are encoded in structured parameters that require further decoding. On Stellar, transactional values are not directly recorded in the standard fields but are embedded within parameters of the function metadata, requiring further decoding for precise financial analysis. In contrast, all detectable transactions on EVM chains: Ethereum, Polygon, Arbitrum, Avalanche, invoke a single function: signedDataExecution. This meta-transaction handler encapsulates pre-authorized operations (e.g., mint, transfer, redeem) within a signed payload, enabling BENJI to enforce off-chain compliance and centralized control over all state changes on EVM chains.
4.3 USDY
We also categorize USDY smart contract functions into operational buckets by applying heuristic string matching on the lowercase decoded function names, aiming to infer their functional semantics. Table 4 describes this mapping, specifically for the USDY token’s diverse DeFi interactions. Given the wide variety of protocols and contract architectures interacting with USDY across multiple chains, the mapping produced aggregation based on the functional semantics. For example, swap captures all decentralized exchange operations regardless of DEX type, while lending encompasses a range of credit-related actions such as borrow, repay, or collateral management.
Figure 2: Function-level distribution of USDY transaction volume vs. frequency across chains. Each point represents a specific $(f,c)$ pair (function $f$ on chain $c$), with marker shape encoding functional category (e.g., swap, lending, execution) and colour denoting blockchain. Both axes use logarithmic scales. This figure reveals how USDY’s cross-chain activity exhibits distinct clusters of functional usage, highlighting protocol specialization (e.g., high-frequency swaps on Mantle vs. large-value mint and burn operations on Ethereum).
Table 4: Functional buckets for USDY contract activity. Buckets are derived by substring pattern matching in decoded function names. These categories reflect common DeFi roles in cross-chain real-world asset token ecosystems.
Function Bucket Name Match (substring) Description
swap swap, unoswap, swaptoken Token swaps via routers or aggregators
liquidity add_liquidity, removeliquidity Liquidity pool provisioning or withdrawal
lending lend, borrow, repay, loan, collateral Lending, borrowing, collateral adjustment
transfer transfer, transfertoken, safetransfer Standard token transfers between addresses
bridge bridge, startbridge, swapandstartbridge Cross-chain bridging of tokens
mint mint Token issuance, typically from off-chain trigger
burn burn Token removal or redemption
rewards claim, harvest, reward, collect Claiming rewards or accrued yield
governance vote, governance Governance interactions (e.g., DAO voting)
execution executemeta, execute, exectransaction, delegatecall, call, multicall General-purpose call wrappers and execution shells
approval approve, permit Token approvals or permission signatures
configuration register, set_, init, config Administrative or protocol configuration actions
unknown (none matched) Uncategorized or obscure logic
Figure 2 visualizes the cross-chain function-level distribution of USDY transactions by plotting total token value against transaction count for each (function, chain) pair. The on-chain data for USDY reveals a mix of institutional-sized wallets and numerous retail holders. The function-level transaction data for USDY reveals a dual-structure usage pattern indicative of both institutional and retail engagement across multiple blockchain environments. Despite USDY’s positioning as a regulated, yield-bearing stablecoin offered under Regulation S [29] to non-U.S. investors, on-chain evidence suggests that the asset has achieved a relatively broad distribution: over 10,000 unique addresses hold USDY across supported chains, suggesting significant uptake by smaller-scale retail participants. However, transaction size and function type diverge markedly between user classes. Large-value, low-frequency operations—such as mint, burn, and high-value execution calls are primarily concentrated on Ethereum mainnet, consistent with institutional-scale issuance and redemption flows. These interactions exhibit transaction sizes in the hundreds of thousands to millions of USDY, aligning with primary market activity and custody-level fund management. By contrast, Layer-2 networks such as Arbitrum and Mantle exhibit distinct transactional profiles characterized by higher frequency, lower median value, and increased heterogeneity in function types (e.g., swap, transfer, and permissioned execution). These patterns are congruent with DeFi-native retail usage, enabled by the lower transaction costs and faster finality on L2s. The structural constraint, though conducive to regulatory compliance, inherently limits composability with standard ERC-20 interfaces and precludes fully permissionless usage; further bifurcating institutional administrative flows from retail DeFi interactions across different layers of the chain stack.
5 Address Role Predictive Modeling
In this section, we focus on address role predictive modeling within the RWA transactions, with the goal of inferring the financial roles of addresses, i.e., not their identity per se, but the economic or operational behaviour they instantiate on-chain, such as traders, execution bots, treasury managers protocol contracts. Even at this level of approximation, functional role inference enables us to potentially contextualize address behaviour, differentiate structural actors in token flows, and support future analysis of token circulation, institutional activity.
Such financial roles can, in part, be derived from existing address labels where available from online resources. However, Labeling on-chain addresses remains a manual and epistemically incomplete task, currently reliant on community-curated, e.g., Dune community-based crypto analytics  [^4] https://dune.com/, or industry-maintained ontologies, e.g., Arkham  [^5] https://intel.arkm.com/, Ethplorer  [^6] https://ethplorer.io/zh/ that emerge retroactively through crowd-sourced investigation, custodial disclosures, or industrial exogenous tagging. These labeling processes are temporally cumulative, i.e., they evolve incrementally over time, contingent on the visibility and social salience of individual addresses. As a result, there exists a persistent gap in functional address understanding at the time of analysis, which necessitates predictive modeling frameworks capable of approximating behavioural roles with minimal supervision. Such automation and approximation provide a practical reference for analysts, partially reducing the cost for manual investigation by offering behaviourally grounded priors that can guide interpretation for manual labeling.
Concretely, we advocate for the inference of financial behaviourally grounded roles: financial controllers treasuries, execution (arbitrage) bots, and retail traders, as proxy labels. In this schema, treasuries refer to multisignature-controlled addresses or custodial vaults (e.g., Gnosis Safe contracts, DAO-managed wallets) that act as long-term capital reserves, execute protocol expenditures, or manage liquidity across chains. The category of execution bots encompasses addresses engaged in high-frequency, low-latency, and often adversarial strategies such as sandwich attacks, MEV extraction, flashloan arbitrage, and Flashbots relaying, whose behavioural signatures deviate from standard market participation and suggest automation and profit maximization. In contrast, retail traders include externally owned accounts (EOAs) that interact primarily with DEXs, NFT marketplaces, or aggregator routers. Such modeling serves both to reduce annotation overhead and potentially enable functional interpretability in transaction graph-level analysis.
5.1 Samples and labels
We aggregate transactions from three representative tokenized real-world assets (RWAs): USDY (Ethereum), BENJI (excluding stellar), and BUIDL. These RWS U.S treasury tokens were selected due to their active cross-chain circulation, institutional provenance, and higher visibility and density of human-readable annotations within community-maintained labeling available sources, making them suitable for semi-supervised role inference.
We extract the naming tags for each address using Dune SQL from the metadata of Dune labels.addresses table  [^7] https://dune.com/data/labels.addresses: an open, curated repository of community-submitted and platform-extracted multi-chain address labels. Each entry includes fields such as name, address, blockchian, source, contributor, etc. We use the name field in the table, which is the naming tags, then apply a set of regular-expression-based rules over the name field to assign coarse-grained functional roles. These rules approximate the aforementioned financial behavioural classes. Table 5 summarizes the pattern-matching logic used to infer each role:
Table 5: Regular expression rules used for coarse address role labeling.
Class Regex
Trader dex trader, aggregator trader, nft trader, daily trader, number of DEXs traded
Bot Sandwich Attacker, Arbitrage, MEV, Flashloan, Flashbots
Treasury Safe, Gnosis Safe, Multisig, DAO Treasury, Vault, Zerion Multisig
Other No match with above patterns
In total, the sampled dataset contains 10,055 transactions between 2023-09-18 and 2025-04-23, covering 815 unique addresses observed as senders or recipients. Using this labeling scheme, we label the address set into four classes: 520 Trader, 33 Bot, 44 Treasury, and 218 Other. These labels serve as training targets for our predictive modeling pipeline, allowing us to benchmark role inference performance under weak supervision.
5.2 Models
RWA transaction graphs exhibit latent financial hierarchy: central actor nodes (e.g., treasuries, issuers, custodians) initiate and coordinate flows, while peripheral nodes (e.g., traders, bots) interact sparsely, locally, or opportunistically with the core. The negative curvature of hyperbolic space provides a proper inductive bias for embedding such structure, enabling compact representations that naturally separate high-degree core nodes near the origin from low-degree boundary nodes along exponentially expanding geodesics. Hence, nodes residing at small hyperbolic radius embody high-hierarchy, systemically central nodes (close to coordinate origin) possibly multisig treasuries, protocol routers, and liquidity hubs whereas larger radius nodes reflect lower-hierarchy participants across the RWA transaction graph. A formal derivation showing that hyperbolic radius encodes hierarchical depth in tree-like graphs where the higher-hierarchy is closer to the originate is provided in Appendix 7, and an illustrative Figure 3 is presented.
Therefore, we propose a hyperbolic node-level representation learning method combined with a feedforward neural networks that integrates transactional features, metadata-driven features including Liquidity-to-Average Ratio (LAR), and hyperbolic (Poincaré) geometry embeddings. The pipeline first encodes latent hierarchy via Poincaré distance-based optimization, then augments each node with hierarchical depth statistics and optional topological features, which are processed by a neural network classifier to infer address roles.
Figure 3: Illustrative example of a rendering of a blockchain transaction graph and its corresponding Poincaré disk embedding. Nodes emulate a three-tier addresses: the red address anchors global liquidity, the green addresses relay funds and manage liquidity pools, and a ring of peripheral trader blue addresses engages sporadically with the core. Embedding the same adjacency structure in the negatively curved Poincaré model separates tiers by geodesic radius.
5.2.1 Poincaré Node Rrepresentation Leanring
Each data point in our framework corresponds to a node $v\in V$ within a token transaction graph $G=(V,E)$, where nodes represent blockchain addresses and directed edges $(u,v)\in E$ denote value transfers. The objective is to learn a high dimensional representation (embedding vector) $\mathbf{z}_{v}\in\mathbb{B}^{d}$ for each address $v$, where $\mathbb{B}^{d}$ is the $d$-dimensional Poincaré ball: a Riemannian manifold of constant negative curvature:
$\mathbb{B}^{d}=\left\{\mathbf{z}\in\mathbb{R}^{d}\;:\;\|\mathbf{z}\|<1\right\}.$
The manifold is equipped with a Riemannian metric that scales the Euclidean inner product via a position-dependent conformal factor:
$g_{\mathbf{z}}=\lambda_{\mathbf{z}}^{2}g^{\mathrm{E}},\qquad\lambda_{\mathbf{z}}=\frac{2}{1-\|\mathbf{z}\|^{2}},$
where $g^{\mathrm{E}}$ denotes the standard Euclidean metric tensor and $\lambda_{\mathbf{z}}$ diverges as $\|\mathbf{z}\|\to 1$, i.e., as points approach the boundary of the ball. The induced geodesic distance between any two points $\mathbf{z}_{u},\mathbf{z}_{v}\in\mathbb{B}^{d}$ is defined by:
$d_{\mathbb{B}}(\mathbf{z}_{u},\mathbf{z}_{v})=\operatorname{arcosh}\left(1+\frac{2\|\mathbf{z}_{u}-\mathbf{z}_{v}\|^{2}}{(1-\|\mathbf{z}_{u}\|^{2})(1-\|\mathbf{z}_{v}\|^{2})}\right),$
[…]
$\mathbf{z}_{u}\oplus\mathbf{z}_{v}=\frac{(1+2\langle\mathbf{z}_{u},\mathbf{z}_{v}\rangle+\|\mathbf{z}_{v}\|^{2})\mathbf{z}_{u}+(1-\|\mathbf{z}_{u}\|^{2})\mathbf{z}_{v}}{1+2\langle\mathbf{z}_{u},\mathbf{z}_{v}\rangle+\|\mathbf{z}_{u}\|^{2}\|\mathbf{z}_{v}\|^{2}}.$
Each update step is followed by projection back onto the open ball to ensure validity:
$\operatorname{proj}(\mathbf{z})=\mathbf{z}\cdot\min\left(1,\frac{1-\varepsilon}{\|\mathbf{z}\|+\varepsilon}\right),$
with $\varepsilon>0$ for numerical stability near the boundary.
We train embeddings via Riemannian stochastic gradient descent (RSGD) over a contrastive objective. Given a positive edge $(i,j^{+})\in E$ and a negative sample $j^{-}\sim\operatorname{Unif}(V)$, we minimize the hinge-based loss:
$\mathcal{L}_{\mathrm{c}}=\frac{1}{|E|}\sum_{(i,j^{+})\in E}\left[d_{\mathbb{B}}(\mathbf{z}_{i},\mathbf{z}_{j^{+}})-d_{\mathbb{B}}(\mathbf{z}_{i},\mathbf{z}_{j^{-}})+\gamma\right]_{+},$
where $\gamma>0$ is a margin hyperparameter and $[\cdot]_{+}=\max(0,\cdot)$ denotes the hinge operator. To align learned radii with graph-theoretic centrality, we regularize embedding norms against normalized degree:
$\mathcal{L}_{\mathrm{r}}=\frac{1}{|V|}\sum_{v\in V}\left(\|\mathbf{z}_{v}\|-\left(1-\frac{\deg(v)}{\max_{u\in V}\deg(u)}\right)\right)^{2},$
encouraging high-degree nodes to concentrate near the center and peripheral actors to occupy the hyperbolic fringe. The total training objective combines contrastive and curvature-alignment terms:
$\mathcal{L}=\mathcal{L}_{\mathrm{c}}+\beta\cdot\mathcal{L}_{\mathrm{r}},$
where $\beta$ (default 0.1) controls the strength of radial regularization. This objective is minimized over certain epochs using intrinsic gradients computed in the Riemannian manifold, followed by Möbius updates and projection. After optimization, each node $v\in V$ is assigned a $\mathbf{z}_{v}\in\mathbb{B}^{64}$, a 64-dimensional hyperbolic embedding that encodes its topological position and connectivity within the latent hierarchy of the transaction graph.
Refinement via Liquidity-to-Average Ratio (LAR) While Poincaré embeddings encode topological position in the latent transactional hierarchy, other features related to token flow transferred within transactions, i.e., the temporal liquidity behaviour that distinguishes roles such as passive holders, programmatic bots, or asymmetric treasuries. To incorporate such information, we define the following over nodes computed on transaction flows.
Definition 5.1 (Liquidity-to-Average Ratio (LAR)).
Let $(u,v)\in E$ denote a directed transaction edge, and let $\mu_{uv}$ and $\sigma_{uv}$ denote the mean and standard deviation of transfer values from $u$ to $v$ within a time window $[t,t+\Delta]$. Let $\texttt{in}(v)$ and $\texttt{out}(v)$ denote the total inflow and outflow of $v$ over the same interval. Then the Liquidity-to-Average Ratio (LAR) for $(u,v)$ is defined as:
$\mathrm{LAR}_{u\to v}=\frac{\sigma_{uv}}{\mu_{uv}+\epsilon}\cdot\left(1+\frac{\sum\texttt{in}(v)}{\sum\texttt{out}(v)+\epsilon}\right),$
where $\epsilon>0$ is a smoothing constant to ensure numerical stability.
$\mathrm{LAR}_{u\to v}$ captures local volatility and directional imbalance: the first term reflects normalized transaction variance; the second penalizes outflow-dominant behaviour. Nodes with high incoming volume and irregular flow patterns exhibit elevated LAR values. To integrate this signal into the hyperbolic embedding space, we aggregate edge-level LAR values into node-level weights. Let $\log(\texttt{val}_{i})$ denote the log-transformed total value received by node $i$, and $\log(\texttt{LAR}_{i})$ the log of its average incident LAR. We compute z-scored forms:
$z^{\mathrm{val}}_{i}=\frac{\log(\texttt{val}_{i})-\mu_{\log(\texttt{val})}}{\sigma_{\log(\texttt{val})}},\quad z^{\mathrm{lar}}_{i}=\frac{\log(\texttt{LAR}_{i})-\mu_{\log(\texttt{LAR})}}{\sigma_{\log(\texttt{LAR})}},$
and define node trust as the sigmoid of their difference: $\tau_{i}=\sigma(z^{\mathrm{val}}_{i}-z^{\mathrm{lar}}_{i})$.
These trust weights modulate a refinement step in hyperbolic space. For each node $i$, let $\mathcal{N}(i)$ denote its neighbours in $G$. Define the trust-weighted tangent update:
$\mathbf{t}_{i}=\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\cdot\log_{0}(\mathbf{z}_{j}),\quad\alpha_{ij}=\frac{\tau_{j}}{\sum_{k\in\mathcal{N}(i)}\tau_{k}},$
where $\log_{0}(\cdot)$ is the logarithmic map at the origin. The new embedding is then updated via Möbius exponential map:
$\mathbf{z}_{i}\leftarrow\operatorname{proj}\left(\tanh\left(\frac{\|\mathbf{t}_{i}\|}{2}\right)\cdot\frac{\mathbf{t}_{i}}{\|\mathbf{t}_{i}\|+\delta}\right),$
with $\delta>0$ ensuring stability near $\|\mathbf{t}_{i}\|=0$. We apply this refinement for $R=3$ steps, with early stopping if embeddings converge.
{observation}
Nodes with low liquidity and irregular behaviour are downweighted during smoothing, preserving geometric sharpness for high-confidence addresses (e.g., well-funded treasuries), while dampening noise from sparsely observed actors.
5.2.2 Hierarchical Radius-based Features
To incorporate latent hierarchy from hyperbolic space, we associate each node $v\in V$ with a Poincaré embedding $\mathbf{z}_{v}\in\mathbb{R}^{d}$ satisfying $\|\mathbf{z}_{v}\|<1$. The node’s hierarchical depth is defined via its hyperbolic radius:
$r_{v}=2\cdot\tanh^{-1}(\|\mathbf{z}_{v}\|),$
which reflects the geodesic distance from the origin and serves as a continuous proxy for depth in the hierarchy, i.e., nodes closer to the origin are higher in the hierarchy, while those near the boundary are deeper.
To capture the local structure around node $v$, we define its $k$-hop neighbourhood $\mathcal{N}_{k}(v)\subseteq V$ (excluding $v$) and collect neighbour radii $\{r_{u}:u\in\mathcal{N}_{k}(v)\}$. From this, we compute an 11-dimensional hierarchical feature vector $\mathbf{h}_{v}\in\mathbb{R}^{11}$ comprising the following components:
$r_{v}$ Absolute depth (node’s self radius)
$\mu$ Mean of neighbour radii: $\mu=\frac{1}{|\mathcal{N}_{k}(v)|}\sum\limits_{u\in\mathcal{N}_{k}(v)}r_{u}$
$\sigma$ Standard deviation of neighbour radii: $\sigma=\sqrt{\frac{1}{|\mathcal{N}_{k}(v)|}\sum\limits_{u\in\mathcal{N}_{k}(v)}(r_{u}-\mu)^{2}}$
$\alpha$ Fraction of neighbours deeper than $v$: $\alpha=\frac{1}{|\mathcal{N}_{k}(v)|}\sum\limits_{u}\mathbb{I}(r_{u}>r_{v})$
$\beta$ Fraction of neighbours shallower than $v$: $\beta=\frac{1}{|\mathcal{N}_{k}(v)|}\sum\limits_{u}\mathbb{I}(r_{u}<r_{v})$
$\delta$ Minimum relative depth: $\delta=\min\limits_{u\in\mathcal{N}_{k}(v)}(r_{u}-r_{v})$
$\Delta$ Maximum relative depth: $\Delta=\max\limits_{u\in\mathcal{N}_{k}(v)}(r_{u}-r_{v})$
$\mathbf{b}_{v}$ 4-bin histogram of relative radii $(r_{u}-r_{v})$ over range $[-1,1]$
Altogether, the hierarchical feature vector is given by:
$\mathbf{h}_{v}=[r_{v},\mu,\sigma,\alpha,\beta,\delta,\Delta\;|\;\mathbf{b}_{v}\in\mathbb{R}^{4}]\in\mathbb{R}^{11},$
providing a compact representation that summarizes both the absolute radial position of node $v$ and the statistical distribution of depths in its local hyperbolic neighbourhood.
5.2.3 Neural Network Inferencer
After obtaining the Poincaré position embedding, and hierarchical radius-based features, we train a supervised neural classifier to infer coarse address roles. Each node $v\in V$ is represented by a concatenated input vectors:
  • $\mathbf{z}_{v}\in\mathbb{R}^{d}$ (constrained to $\mathbb{B}^{d}$): Poincaré embedding (cf. Section 5.2.1);
  • $\mathbf{h}_{v}\in\mathbb{R}^{11}$: hierarchical features computed from local radius-based statistics (cf. Section 5.2.2);
  • $\mathbf{r}_{v}\in\mathbb{R}^{k}$: vector of topology-aware node features derived from truncated random walks over local neighbourhoods (as in DeepWalk embeddings [16]).
These components are concatenated into a unified feature vector:
$\mathbf{x}_{v}=\left[\mathbf{z}_{v}\,\middle\|\,\mathbf{r}_{v}\,\middle\|\,\mathbf{h}_{v}\right]\in\mathbb{R}^{d+k+11},$
where $d$ is the Poincaré embedding dimension, $k$ is the raw feature dimension. The neural architecture consists of a two-layer multilayer perceptron (MLP) with batch normalization, ReLU activations, and dropout regularization. Letting $\mathrm{MLP}:\mathbb{R}^{d+k+11}\rightarrow\mathbb{R}^{C}$ denote the classifier, the unnormalized logits for node $v$ are:
$\hat{\mathbf{y}}_{v}=\mathrm{MLP}(\mathbf{x}_{v})=\mathbf{W}_{2}\cdot\mathrm{ReLU}\left(\mathbf{W}_{1}\cdot\mathbf{x}_{v}+\mathbf{b}_{1}\right)+\mathbf{b}_{2},$
where $C$ is the number of target classes. Training is supervised using the standard cross-entropy loss over weak role labels $\{y_{v}\}_{v\in\mathcal{D}_{\mathrm{train}}}$:
$\mathcal{L}_{\mathrm{clf}}=-\sum_{v\in\mathcal{D}_{\mathrm{train}}}\log\frac{\exp(\hat{y}_{v}^{(y_{v})})}{\sum_{c=1}^{C}\exp(\hat{y}_{v}^{(c)})}.$
This loss is minimized when the model assigns high probability to the correct class. The ground-truth label $y_{v}$ selects the corresponding logit $\hat{y}_{v}^{(y_{v})}$ from the model’s output $\hat{\mathbf{y}}_{v}$, and the loss penalizes the model when this logit does not dominate the softmax distribution, i.e., when the predicted probability for class $y_{v}$ is low.
5.2.4 Experiment Result
We evaluate our role classifier using a stratified train/test split with an 80/20 ratio over the set of labeled addresses with the four coarse-grained role classes (cf. Section 5.1). The model is optimized using the AdamW [13] optimizer with a learning rate of $10^{-3}$ and a weight decay of $0.2$. Training is conducted for up to 2k epochs with early stopping based on macro-F1 score on the validation set, using a patience threshold of 10 epochs.
The best-performing model checkpoint (by macro-F1) is selected for final evaluation. We conduct ablation study using with the input features of Poincaré embeddings $\mathbf{z}_{v}\in\mathbb{R}^{64}$, hierarchical descriptors $\mathbf{h}_{v}\in\mathbb{R}^{11}$ ($k=1$), and topology-aware node embeddings $\mathbf{r}_{v}\in\mathbb{R}^{64}$, with feature subsets in combination under different ablation settings.
We named our model PoincaVec and compare the method against three established node representation baselines: Node2Vec [10], Role2Vec [1], and FeatherNode [19]. We select Role2Vec, and FeatherNode as baselines for comparison due to their established efficacy in modeling blockchain transaction graphs, as evidenced in prior evaluations [8, 6]. We adopt the same hyperparameter settings used in prior blockchain transaction address representation learning works [17, 6] , setting the context size $10$, embedding dimension $64$, walk length $5$, number of walks per node $10$ for both Node2Vec and Role2Vec, while FeatherNode is configured SVD iterations 20. The baseline Node2Vec [10] captures basic neighbour homophily-based proximity via biased random walks, while Role2Vec [1] models structural equivalence by clustering nodes based on topological statistics (e.g., degree, triangle count, k-core number) into roles (types) and learning role-level embeddings where the nodes share the same role obtain the same embedding vector. FeatherNode [19] incorporates spectral and node-level distributional features from attributed graphs by applying Fourier transforms of empirical distributions to neighbourhood feature aggregations. These aggregations are subsequently compressed using truncated Singular Value Decomposition (SVD), yielding node embeddings that are invariant to permutation and sensitive to local graph structure.
Table 6: Role classification performance on the RWA dataset. H: hierarchical radius-based features; T: topology-aware node features.
Model Precision Recall F1 Accuracy
Node2Vec 0.668 0.687 0.654 0.687
Role2Vec 0.659 0.681 0.662 0.681
FeatherNode 0.407 0.638 0.497 0.638
PoincaVec (w/o H, w/o T) 0.692 0.712 0.694 0.712
PoincaVec (w/ H, w/o T) 0.710 0.706 0.684 0.706
PoincaVec (w/ H, w/ T) 0.757 0.748 0.726 0.748
Table 6 presents the classification performance of PoincaVec and several baseline embedding methods on the RWA dataset. Among the baselines, Node2Vec and Role2Vec provide modest F1 scores of 0.654 and 0.662, respectively, while FeatherNode underperforms 0.497, likely due to the high sparsity and low feature homogeneity in financial transaction graphs.
Without hierarchical or walk-based features, PoincaVec (w/o H, w/o T), which uses only Poincaré embeddings, outperforms all baselines with an F1 of 0.694. Concatenating hierarchical radius features (w/ H, w/o T) maintains competitive performance, while integrating topology-aware node vectors derived from truncated random walks (w/ H, w/ T) leads to a result F1 of 0.726. These results demonstrate that curvature-aware embeddings, when enriched with role-informed and structural signals, provide competitive representational capacity for role inference in blockchain transaction graphs.
Evaluation on External Blockchain Transaction Datasets To further evaluate the generalization capacity of our PoincaVec architecture beyond the RWA role classification task, we test it on multiple publicly available blockchain transaction graph datasets with labels. These datasets include addresses labeled as fraudulent or scam-related, whose topological irregularities render them particularly amenable to detection using curvature-aware embeddings and hierarchical structural descriptors. Given that fraud or anomalous addresses often lie structurally at the periphery or occupy non-homophilic positions in the graph, the curvature-aware inductive bias of PoincaVec may help reveal subtle hierarchical distinctions absent in flat Euclidean spaces.
Ethereum Transaction [31]
An Ethereum transaction transaction dataset constructed from historical blockchain records, containing 2,973,489 addresses and 13,551,303 edges. A total of 1,165 addresses are labeled as illicit based on curated forensic data, while the rest remain unlabeled. This dataset has an average node degree of 4.56.
AscendEXHacker [30]
A subgraph (network) extracted via Etherscan’s Heist label, focusing specifically on the 2021 AscendEX exchange exploit. It contains 6,642 addresses across 29,074 transactions. Ground-truth annotations label 84 addresses as direct heist participants, while 638 addresses are identified as DEXs or Uniswap-related service accounts.
PlusTokenPonzi [30]
A transaction dataset linked to the PlusToken Ponzi scheme, one of the largest crypto scams to date. It comprises 34,521 unique accounts and 58,049 transactions, of which 30,782 addresses are explicitly identified as scam participants.
Ethereum Classic Dataset [2]
A labeled transaction network from the Ethereum Classic (ETC) network, compiled using EtherscamDB scam reports. The dataset comprises 73,034 nodes and 71,250 edges, with 2,357 addresses labeled as scammers based on crowdsourced and externally verified annotations.
Table 7: Role classification performance across four blockchain transaction datasets.
Ethereum
Model Precision Recall F1 Accuracy
Node2Vec 0.922 0.918 0.918 0.918
Role2Vec 0.923 0.923 0.923 0.923
FeatherNode 0.922 0.916 0.916 0.916
PoincaVec (w/ H, w/o T) 0.901 0.897 0.897 0.896
PoincaVec (w/o H, w/ T) 0.936 0.936 0.936 0.936
PoincaVec (w/ H, w/ T) 0.941 0.940 0.940 0.940
AscendEXHacker
Node2Vec 0.889 0.719 0.780 0.991
Role2Vec 0.829 0.675 0.728 0.989
FeatherNode 0.694 0.558 0.587 0.986
PoincaVec (w/ H, w/o T) 0.935 0.779 0.840 0.993
PoincaVec (w/o H, w/ T) 0.925 0.739 0.806 0.992
PoincaVec (w/ H, w/ T) 0.897 0.735 0.794 0.992
PlusTokenPonzi
Node2Vec 0.996 0.990 0.993 0.996
Role2Vec 0.993 0.996 0.994 0.996
FeatherNode 0.996 0.992 0.994 0.996
PoincaVec (w/ H, w/o T) 0.995 0.979 0.987 0.992
PoincaVec (w/o H, w/ T) 0.997 0.991 0.994 0.996
PoincaVec (w/ H, w/ T) 0.996 0.986 0.991 0.994
Ethereum Classic (ETC)
Node2Vec 0.910 0.898 0.899 0.899
Role2Vec 0.888 0.882 0.882 0.882
FeatherNode 0.927 0.926 0.926 0.926
PoincaVec (w/ H, w/o T) 0.910 0.898 0.899 0.900
PoincaVec (w/o H, w/ T) 0.921 0.913 0.914 0.914
PoincaVec (w/ H, w/ T) 0.949 0.943 0.945 0.946
The results in Table 7 compare the performance of benchmark methods and our proposed PoincaVec variants across four blockchain transaction graph datasets with varying structural properties and label sparsity. Across all datasets, the full PoincaVec pipeline incorporating both hierarchical radius-based features (H) and topology-aware walk embeddings (T) achieves the highest or near-highest F1 scores, demonstrating strong generalization to fraud node (address) classification tasks.
For the Ethereum dataset featuring a large, sparsely labeled transaction network, PoincaVec (w/ H, w/ T) achieves an F1 of 0.940, outperforming proximity-based baselines such as Node2Vec and structurally driven Role2Vec. The AscendEXHacker dataset exhibits an imbalanced role distribution, where only a small fraction of addresses are tagged as direct heist participants, while the majority are exchange-facing services such as DEX routers (e.g., Uniswap). In this setting, hierarchical radius-based features (H) offer an inductive prior: heist addresses tend to lie peripherally with sparse connectivity, while services cluster near radius enabling PoincaVec (w/ H, w/o T) to achieve the highest F1 of 0.840, whereas adding walk-based features reduces performance to 0.794. Notably, PoincaVec obtained an F1 of 0.945 on the Ethereum Classic (ETC) dataset, demonstrating its capacity to disentangle abnormal node roles. These results demonstrate the performance of embedding blockchain transaction graphs in hyperbolic space, where latent hierarchies and peripheral node behaviours, which are common in fraud and scam context, can be captured more compactly in our proposed method.
6 Conclusion
This study presents the first transaction-level analysis of tokenized U.S. Treasuries across multiple blockchain networks. By decoding contract interactions and profiling function usage, we identify dominant operational roles and reveal distinct patterns between institutional and retail participants. Our cross-chain comparison highlights how token design and network context shape transactional behaviour. Furthermore, we introduce a hyperbolic embedding model that infers address roles. These findings provide a quantitative account of transaction behaviours across tokenized U.S. Treasuries, detailing how operational functions and participant roles emerge across chains and transactional contexts. The proposed Poincaré-embedding-based model enables role identification and achieves improved performance both on our RWA dataset and in generalizing to external classification and anomaly detection tasks.
7 Appendix: Hyperbolic Radius as a Proxy for Latent Hierarchy
Let $\mathbb{T}_{k}$ denote a $k$-ary tree, and consider a mapping $\phi:\mathbb{T}_{k}\rightarrow\mathbb{B}^{d}$ from a tree into the $d$-dimensional Poincaré ball $\mathbb{B}^{d}=\{\mathbf{z}\in\mathbb{R}^{d}\mid\|\mathbf{z}\|<1\}$.
Let $u$ be a node at depth $h$ in the tree (i.e., $h$ hops from the root). Suppose $\phi$ maps nodes along a fixed geodesic direction such that the hyperbolic distance from the origin encodes depth:
$r_{h}:=\|\phi(u)\|=\tanh\left(\frac{h\cdot\ell}{2}\right),$
where $\ell>0$ is the fixed geodesic step length in the hyperbolic metric. Then:
Lemma 7.1.
The hyperbolic radius $r_{h}$ is a strictly increasing function of tree depth $h$.
Proof 7.2.
Since $\tanh(x)$ is a strictly increasing function for $x>0$, and $h\mapsto h\ell/2$ is linear and positive for $h\geq 0$, we have:
$\frac{d}{dh}r_{h}=\frac{d}{dh}\tanh\left(\frac{h\ell}{2}\right)=\frac{\ell}{2}\cdot\text{sech}^{2}\left(\frac{h\ell}{2}\right)>0.$
Therefore, $r_{h}$ increases monotonically with $h$.
Corollary 7.3.
Under tree-consistent Poincaré embeddings, node radius $\|\mathbf{z}_{v}\|$ can be interpreted as a continuous proxy for latent hierarchical depth.